OpenStudy (anonymous):

PLZZZ HELP!!!!! I WILL MEDLE AND FAN Shonda made a model of a solid ice cream cone with a spherical scoop of ice cream on top, made out of clay. Which formulas are needed to calculate how much clay she used? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.) A. B. C. D.

OpenStudy (anonymous):

a: V=1/3*3.14*R2H

OpenStudy (anonymous):

B:V=*3.14*R2H

OpenStudy (anonymous):

C:V=LWH

OpenStudy (anonymous):

D:V=4/3*3.14*R3

OpenStudy (anonymous):

i think its a

OpenStudy (nerdgirl):

Take look at this and let me know if it helps :) http://openstudy.com/study#/updates/54f20c99e4b09f47f4e3cad9 and no, it's not a

OpenStudy (mathstudent55):

You need the formulas for the volume of a sphere and volume of a cone.

OpenStudy (anonymous):

oh so is it d?

OpenStudy (mathstudent55):

\(V_{cone} = \dfrac{1}{3} \pi r^2 h\) \(V_{sphere} = \dfrac{4}{3} \pi r^3\)

OpenStudy (mathstudent55):

a. is the volume of a cone b. is the volume of a cylinder c. is the volume of a rectangular prism d. is the volume of a sphere You need both the cone and the sphere.

OpenStudy (anonymous):

ohh so its a and d

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

wait but i was right on both answers tho wasent i?

OpenStudy (mathstudent55):

Yes, but it's not just a. or just d. It's both a. and d. You need to pick all formulas that are needed, not just one of them.

OpenStudy (mathstudent55):

Both formulas you picked were correct. Answer is: a. and d.

OpenStudy (anonymous):

ohh ok thanks can you help me with 4 more questions?

OpenStudy (anonymous):

@mathstudent55 or @mathsciencehistory

OpenStudy (anonymous):

Two spherical balloons are filled with water. The first balloon has a radius of 3 cm, and the second has a radius of 6 cm. How much more water is in the larger balloon than in the smaller balloon? Use 3.14 to approximate pi, and express your answer in the hundredths. cm3

OpenStudy (anonymous):

_______cm3